OpenStudy (anonymous):

A police officer measured a skid mark to be 150 ft long. The driver said he was going under the speed limit, which was 45 miles per hour. Was the driver correct? How long would the skid mark be if he was indeed going 45 miles per hour when the brakes were applied?

OpenStudy (dumbcow):

what is the deceleration force applied by the brakes?

OpenStudy (anonymous):

That is all it says

OpenStudy (dumbcow):

If d is length of skid mark going 45mph or 66 ft/sec d = 66^2/2k where k is acceleration of the brakes

OpenStudy (anonymous):

ty

OpenStudy (dumbcow):

found a formula that might be useful \[v = \sqrt{30(0.7)d}\] where 0.7 refers to drag coefficient for asphalt and assuming ideal braking efficiency similarly then \[d = \frac{v^{2}}{(30)(0.7)}\] using this for v=45 \[d = 96.4 ft\] driver was obviously not going 45 mph for d=150 \[v = 56.1 mph\]